A particle is projected with a velocity 60m/s at an angle 60 from horizontal then the velocity of the particle when it makes angle 30 with the horizontal is?

Grade:11

3 Answers

Ankur Saha
7 years ago
60* cos60= x* cos30
==> x=60/1.732 = 34.64 m/s
hence the velocity of the particle is 34.64 m/s................................................................................................................................................................................
DR STRANGE
7 years ago
hi,
the horizontal velocity should remain constant!
initially its 60*cos600
let the particle will have velocity x m/s  at angle 300
so 60*cos60= x*cos300
x =\frac{60}{\sqrt{3}} =20\sqrt{3} m/s
DR STRANGE
7 years ago
plz approve if you are sat